### v1

# Kuowen Chen

Abstract
In this paper we consider the problem of approximating Euclidean distances by the infinite integer grid graph. Although the topology of the graph is fixed, we have control over the edge-weight assignment $w:E\to\mathbb{R}_{\geq 0}$, and hope to have grid distances be asymptotically isometric to Euclidean distances, that is:
For all grid points $u,v$, $\operatorname{dist}_{w}(u,v)=(1\pm o(1))\|u-v\|_{2}$.
We give three methods for solving this problem, each attractive in its own way.
  • Our first construction is based on an embedding of the recursive, non-periodic pinwheel tiling of Radin and Conway [Rad94, RS96, CR98] into the integer grid. Distances in the pinwheel graph are asymptotically isometric to Euclidean distances, but no explicit bound on the rate of convergence was known. We prove that the multiplicative distortion of the pinwheel graph is $(1+1/\Theta(\log^{\xi}\log D))$, where $D$ is the Euclidean distance and $\xi=\Theta(1)$. The pinwheel tiling approach is conceptually simple, but can be improved quantitatively.
  • Our second construction is based on a hierarchical arrangement of highways. It is simple, achieving stretch $(1+1/\Theta(D^{1/9}))$, which converges doubly exponentially faster than the pinwheel tiling approach.
[…]
1 Introduction
In this paper we consider a natural geometric problem tangentially related to metric embeddings, spanners, and, in its randomized form, percolation theory. Suppose we wish to approximate Euclidean distances between points on the plane, but with a simple discrete structure: the integer grid graph $\mathsf{Grid}=(\mathbb{Z}\times\mathbb{Z},\{\{u,v\}\mid\|u-v\|_{1}=1\})$. If we consider all edges of $E(\mathsf{Grid})$ to have unit length, then $\mathsf{Grid}$ can be regarded as a $\sqrt{2}$-spanner since for any $(u,v)\in(\mathbb{Z}^{2})^{2}$,
$\|u-v\|_{2}\leq\operatorname{dist}_{\mathsf{Grid}}(u,v)\leq\sqrt{2}\cdot\|u-v\|_{2}.$
Now define $\mathsf{Grid}[w]$ to be $\mathsf{Grid}$ endowed with a non-negative edge-weight assignment $w:E(\mathsf{Grid})\to\mathbb{R}_{\geq 0}$, and let $\operatorname{dist}_{w}$ be the distance function with respect to $w$. We consider the natural question: does there exist a $\mathsf{Grid}[w^{*}]$ that is an asymptotic 1-spanner of the Euclidean plane?
Question 1 (The Squishy Grid Problem).
Does there exist a weight function $w^{*}$ such that for all $u,v\in V(\mathsf{Grid})$ $\operatorname{dist}_{w^{*}}$ is asymptotically Euclidean? That is,
$\operatorname{dist}_{w^{*}}(u,v)=(1\pm o(1))\|u-v\|_{2}.$
If so, we may distinguish various types of convergence:
Polynomial.
$\operatorname{dist}_{w^{*}}(u,v)=\|u-v\|_{2}\pm O(\|u-v\|_{2})^{1-\Omega(1)}$.
Subpolynomial.
$\operatorname{dist}_{w^{*}}(u,v)=\|u-v\|_{2}\pm(\|u-v\|_{2})^{o(1)}$.
Constant.
$\operatorname{dist}_{w^{*}}(u,v)=\|u-v\|_{2}\pm O(1)$.
Before discussing our approach to answering Question 1 we review the history of Question 1 and its connections to percolation theory.
1.1 History of the Problem and Related Results
G. Tardos (personal communication) made us aware of a 1990 book chapter of Pach, Pollack, and Spencer [PPS90] who attributed some version of Question 1 to Paul Erdős. Pach et al. [PPS90] proved that for any fixed $\epsilon>0$ there is a weight function $w[\epsilon]$ such that for all $u,v\in V(\mathsf{Grid})$,
$\|u-v\|_{2}\leq\operatorname{dist}_{w[\epsilon]}(u,v)\leq(1+\epsilon)\|u-v\|_{2}+O(5^{1/\epsilon}),$
which does not resolve Question 1. Borradaile and Eppstein [BE15] considered a more general problem: given a point set $P\in\mathbb{R}^{2}$, compute a weighted planar graph $G=(P\cup S,E)$ with Steiner points $S$ such that $\operatorname{dist}_{G}(u,v)$ $(1+\epsilon)$-approximates the Euclidean distance $\|u-v\|_{2}$. They proved that $|S|=O_{\epsilon,\alpha}(|P|)$ suffices, where $\alpha$ is the sharpest angle in the Delaunay triangulation of $P$. A result of Chang, Krauthgamer, and Tan [CKT22] implies an upper bound of $O_{\epsilon}(|P|\operatorname{polylog}|P|)$, which is slightly superlinear but independent of $\alpha$.
The problem was first posed to us by G. Bodwin, not as a deterministic design problem (Question 1) but as a randomized one. Whenever $\mathscr{D}$ is a distribution over $\mathbb{R}_{\geq 0}$, let $\mathsf{Grid}[\mathscr{D}]$ be the distribution of weighted graphs such that for each $e\in E(\mathsf{Grid})$, $w(e)\sim\mathscr{D}$ is sampled independently from the distribution. Is it possible to find a distribution $\mathscr{D}^{*}$ such that distances in $\mathsf{Grid}[\mathscr{D}^{*}]$ are Euclidean in expectation? In more detail:
Question 2 (Randomized Squishy Grid Problem).
Does there exist a distribution $\mathscr{D}^{*}$ over $\mathbb{R}_{\geq 0}$ such that if $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}^{*}]$ is a randomly weighted graph, for all $u,v\in V(\mathsf{Grid})$,
${\mathbb{E}\/}(\operatorname{dist}_{w}(u,v))=(1\pm o(1))\|u-v\|_{2}.$
The randomized process implicit in Question 2 is actually not new, but dates back to at least a 1965 paper of Hammersley and Welsh [HW65], who called it first passage percolation. They imagined an orchard in which trees were planted on the integer lattice. One tree is initially infected, and the time taken for an infected tree to infect a cardinal neighbor is governed by a distribution $\mathscr{D}$ on $\mathbb{R}_{\geq 0}$. One can then ask: how far does the infection spread by time $t$? and what does the set of infected trees look like?
Many basic questions in first passage percolation theory remain open, and we can quickly summarize the known facts related to Question 2. Let $\mathbf{0}=(0,0)$ be the origin and $e_{\theta}$ be the unit vector with angle $\theta$ degrees. We interpret $ne_{\theta}$ to mean the integer point in $V(\mathsf{Grid})$ nearest to $ne_{\theta}$. The time constant $\mu_{0}(\mathscr{D})$ is such that $\lim_{n\to\infty}\operatorname{dist}_{w}(\mathbf{0},ne_{0})/n=\mu_{0}$ almost surely, which exists if, whenever $w_{1},\ldots,w_{4}\sim\mathscr{D}$ are independently sampled, ${\mathbb{E}\/}(\min\{w_{1},w_{2},w_{3},w_{4}\})<\infty$ [Kes86]. It follows that $0\leq\mu_{0}\leq{\mathbb{E}\/}(w_{1}\sim\mathscr{D})$, with the latter inequality holding with equality only if $w_{1}\sim\mathscr{D}$ is constant almost surely [HW65]. Similarly, the time constants for other angles $\mu_{\theta}(\mathscr{D})=\lim_{n\to\infty}\operatorname{dist}_{w}(\mathbf{0},ne_{\theta})/n$ exist, and collectively define the limiting shape of the balls under distribution $\mathscr{D}$. Let $B(t)=\{u\in\mathbb{Z}^{2}\mid\operatorname{dist}_{w}(\mathbf{0},u)\leq t\}$ be the ball of radius $t$ around the origin. The Cox-Durrett shape theorem [CD81] shows that with probability 1, as $t\to\infty$, $B(t)/t$ tends to a fixed limit shape $\mathcal{B}(\mathscr{D})\subset\mathbb{R}^{2}$. When $\mu_{0}(\mathscr{D})>0$, $\mathcal{B}(\mathscr{D})$ is bounded, convex, and has the same symmetries as $\mathbb{Z}^{2}$, and when $\mu_{0}(\mathscr{D})=0$, $\mathcal{B}(\mathscr{D})$ is $\mathbb{R}^{2}$ itself. See [ADH17] for an extensive survey of first passage percolation theory.
In the context of answering Question 2 we can rescale any non-trivial distribution $\mathscr{D}$ so that its time constant $\mu_{0}(\mathscr{D})=1$, i.e., distances from the origin to points on the $x$- and $y$-axes are asymptotically isometric. In light of the Cox-Durrett theorem, Question 2 asks whether there exists a $\mathscr{D}$ for which $\mathcal{B}(\mathscr{D})$ is the unit $L_{2}$ ball $\{x\mid\|x\|_{2}\leq 1\}$.
Unfortunately, there are no results characterizing $\mathcal{B}(\mathscr{D})$ for any non-trivial distribution $\mathscr{D}$. It is not even known whether there exists $\mathscr{D}$ such that
$\lim_{n\to\infty}\frac{{\mathbb{E}\/}(\operatorname{dist}_{w}(\mathbf{0},ne_{0}))}{n}=\lim_{n\to\infty}\frac{{\mathbb{E}\/}(\operatorname{dist}_{w}(\mathbf{0},ne_{45}))}{n}=1,$ (1)
i.e., $\mathcal{B}(\mathscr{D})$ coincides with the unit $L_{2}$-ball on the eight (inter)cardinal directions. On the other hand, we have solid experimental evidence that $\mathcal{B}(\mathscr{D})$ can get within 1% of the unit $L_{2}$-ball, for certain distributions $\mathscr{D}$. A study of Alm and Deijfen [AD15] looked at various continuous distributions $\mathscr{D}$. When $\mathscr{D}$ is the uniform distribution, the limit shape $\mathcal{B}(\mathscr{D})$ approximates the $L_{2}$-ball to with 4%, whereas when $\mathscr{D}$ is exponential the limit shape is about 1.5% away from the $L_{2}$-ball. The best empirical approximation to the $L_{2}$-ball came from a Fisher distribution, with error less than 1%.
1.2 Results and Findings
We provide two approaches to answering Question 1. and present additional experimental evidence that Question 2 can be answered in the affirmative, using simple discrete distributions.
Our first construction is based on Radin and Conway’s pinwheel tiling [Rad94, RS96, CR98], a conceptually simple tiling that emerges from the observation that a right triangle with proportions $1:2:\sqrt{5}$ can be partitioned into five right triangles with the same proportions. It is known [RS96] that when regarded as a plane graph $G_{\operatorname{PW}}$ with edges weighted according to Euclidean distance, distances in the pinwheel tiling are Euclidean in the limit, that is,
$\lim_{d\to\infty}\;\max_{u,v:\|u-v\|>d}\;\frac{\operatorname{dist}_{G_{\operatorname{PW}}}(u,v)}{\|u-v\|_{2}}=1.$
However the rate of convergence is unknown. We embed the pinwheel tiling into the grid graph, and prove a bound on its convergence, namely that for a constant $\xi=\Theta(1)$,
$\operatorname{dist}_{G_{\operatorname{PW}}}(u,v)=\left(1+O\left(\frac{1}{\log^{\xi}\log\|u-v\|_{2}}\right)\right)\|u-v\|_{2}.$
A natural problem is to optimize the convergence rate of the construction. We give a new, simple construction of a weight function $w$ of the grid that is asymptotically Euclidean, with a polynomial convergence rate.
$\operatorname{dist}_{w}(u,v)=\|u-v\|_{2}+O(\|u-v\|_{2}^{8/9}).$
[…]
Alm and Deijfen’s [AD15] experimental study of first passage percolation selected $\mathscr{D}$ from various continuous distributions such as uniform, exponential, Gamma, and Fisher distributions. For several of these distributions the limit shape $\mathcal{B}(\mathscr{D})$ approximated the $L_{2}$-ball within a few percent, with a Fisher distribution being the best. Our experiments show that very simple distributions with support size 2 or 3 can replicate the accuracy of the continuous distributions [AD15]. For example, the improbable distribution $\mathscr{D}_{2}$:
$\Pr_{w_{0}\sim\mathscr{D}_{2}}\left(w_{0}=\left\{\begin{array}[]{l}0.41401\\ 4.75309\end{array}\right.\right)=\left\{\begin{array}[]{l}0.44273\\ 0.55727\end{array}\right.$
[…]
1.3 Organization
We present the construction based on pinwheel tilings in Section 2, as well as new bounds on the convergence of the stretch of the pinwheel graph. The highway construction is presented in Section 3, having polynomial convergence. We present the experimental findings in Section 4 and conclude with a discussion of several open problems related to Questions 1 and 2 in Section 5.
Sections 2, 3, and 4 are written to be entirely independent. One may read them in any order.
2 A Deterministic Construction Based on Pinwheel Tilings
The “pinwheel” tiling of Radin [Rad94] is an example of a non-periodic tiling using a single tile type (and its reflection). Let $\mathscr{T}_{0},\mathscr{T}_{1},\mathscr{T}_{2},\ldots$ be a series of tilings of ever larger triangular swatches of the plane, and let $\mathscr{T}_{\omega}$ be the tesselation of the plane achieved in the limit. $\mathscr{T}_{0}$ consists of a single right triangle with side lengths $1,2,\sqrt{5}$. In general $\mathscr{T}_{i+1}$ is formed from $\mathscr{T}_{i}$ by taking four additional copies of $\mathscr{T}_{i}$, suitably reflected, rotated, and translated, so that they form a larger triangle with the same $1:2:\sqrt{5}$ proportions. Fig. 1 illustrates the construction of $\mathscr{T}_{2}$ from $\mathscr{T}_{1}$ and $\mathscr{T}_{0}$.
2.1 Pinwheel Tilings
By construction $\mathscr{T}_{\omega}$ is a tiling of the plane using atomic triangles with side lengths $1,2,\sqrt{5}$. Due to the recursive nature of the construction, we can also regard $\mathscr{T}_{\omega}$ as a tiling using $\sqrt{5}^{i},2\sqrt{5}^{i},\sqrt{5}^{i+1}$ triangles, for any integer $i\geq 0$. Observe from Fig. 1 that the boundary of $\mathscr{T}_{i+1}$ is obtained from the boundary $\mathscr{T}_{i}$ by scaling by $\sqrt{5}$, translation, and rotation by $\arctan(1/2)$. We will henceforth define $\gamma=\arctan(1/2)$. As $\gamma/(2\pi)$ is irrational, the orientation of tiles in $\mathscr{T}_{\omega}$ is uniformly distributed in $[0,2\pi)$. Radin and Sadun [RS96] used this fact to prove an isoperimetric property of pinwheel tilings, namely that there are finite subsets of tiles from $\mathscr{T}_{\omega}$ whose area/perimeter2 is arbitrarily close to that of the circle. Suppose we regard $\mathscr{T}_{\omega}$ as a plane graph $G_{\omega}$, whose vertices and edges are the union of the vertices and edges of all atomic triangles. Radin and Sadun [RS96] proved that for $u,v\in V(G_{\omega})$, $\operatorname{dist}_{G_{\omega}}(u,v)=(1+o(1))\|u-v\|_{2}$. Although the multiplicative stretch is 1 in the limit, their proof implies no particular rate of convergence. We prove the following.
(A) (B)
Figure 1: (A) $\mathscr{T}_{1}$, containing $\mathscr{T}_{0}$ in red. (B) $\mathscr{T}_{2}$, containing $\mathscr{T}_{1}$ in red, and $\mathscr{T}_{0}$ in blue.
Theorem 2.1.
Let $G_{\omega}$ be the plane graph of the pinwheel tiling $\mathscr{T}_{\omega}$, whose edges are weighted according to the Euclidean distance between their endpoints. Then for any $u,v\in V(G_{\omega})$,
$\|u-v\|_{2}\leq\operatorname{dist}_{G_{\omega}}(u,v)\leq(1+O(1/(\log\log\|u-v\|_{2})^{\xi}))\cdot\|u-v\|_{2},$
for some $\xi>0$.
2.2 Distribution of Tile Orientations
[…]
Observation 2.2.
Suppose a triangle $T$ of level $x$ has its hypotenuse at angle $\theta$. For $0\leq k\leq x$, $\mathrm{Angles}(T,k)\supseteq\left\{\theta+(2t-k)\gamma\,|\,t\in\{0,\ldots,k\}\right\}$.
Proof.
[…]
$\mathrm{Angles}(A,k-1)$$\supseteq\{\theta-\gamma+(2t-k+1)\gamma\,\,|\,\,t\in 0,\ldots,k-1\}=\{\theta+(2t-k)\gamma\,\,|\,\,t\in 0,\ldots,k-1\}$
$\mathrm{Angles}(B,k-1)$$\supseteq\{\theta+\gamma+(2t-k+1)\gamma\,\,|\,\,t\in 0,\ldots,k-1\}=\{\theta+(2t-k)\gamma\,\,|\,\,t\in 1,\ldots,k\},$
and the union of these sets is exactly $\left\{\theta+(2t-k)\gamma\,|\,t\in\{0,\ldots,k\}\right\}$.
BA
Figure 2: Two triangles within $T$.
Consider some $\lambda\in\mathbb{R}$, and denote by $\{\lambda\}=\lambda-\lfloor\lambda\rfloor$ the fractional part of $\lambda$. It is well known that if $\lambda$ is irrational, the set $S(\lambda,N)=\left\{\{kx\}:1\leq k\leq N\right\}$ becomes uniformly distributed in [0,1)[0,1)[ 0 , 1 ) as $N\to\infty$. Moreover, the rate of convergence is controlled by the irrationality exponent $\mu(\lambda)$ which measures the asymptotic quality of rational approximations to $\lambda$.
For a real number $\lambda$, $\mu(\lambda)$ is defined as the smallest positive real number such that there exists a constant $c(\lambda,\mu)$ for which
$0<\left|\lambda-\frac{p}{q}\right|<\frac{c(\lambda,\mu)}{q^{\mu}}$
has no solutions for rational $p/q$. A finite irrationality exponent implies that a number does not have rational approximations that are “too good.” Moreover, if $\mu(\lambda)$ is finite, $S(\lambda,n)$ is an $\varepsilon$-cover of [0,1)[0,1)[ 0 , 1 ) for $N=O(\varepsilon^{1-\mu(\lambda)})$. The following corollary follows by applying this fact to 2.2.
Corollary 2.3.
[…]
Before bounding the irrationality exponent $\mu(\gamma/\pi)$, we quickly review some terminology related to the algebraic numbers $\overline{\mathbb{Q}}$. For $\alpha\in\overline{\mathbb{Q}}$, its minimal polynomial is the unique polynomial $P\in\mathbb{Z}[x]$ of lowest degree with relatively prime coefficients such that $P(\alpha)=0$. We say the degree of $\alpha$ is the degree of $P$, while the height of $\alpha$ is the absolute value over coefficients of $P$.
Both $\gamma$ and $\pi$ can be expressed in the form $\beta\ln\alpha$ where $\alpha\in\overline{\mathbb{Q}}$ and $\beta\in\mathbb{Q}(i)$. Since $e^{i\gamma}=\frac{2+i}{\sqrt{5}}$, $\gamma=-i\ln{(\frac{2+i}{\sqrt{5}})}$. Similarly, $\pi$ can be written as $-i\ln{(-1)}$. A great deal of work in the mid-20th century yielded various lower bounds on linear forms in logarithms of algebraic numbers. A history of the problem up to 1976 can be found in [Bak76]. For our purposes, we choose a simple bound due to Baker.
Theorem 2.4 (Baker [Bak76]).
Suppose for $n\geq 1$ we have algebraic numbers $\alpha_{1},\ldots,\alpha_{n},\beta_{1},\dots,\beta_{n}\in\overline{\mathbb{Q}}\setminus 0$. If the logarithms $\ln{\alpha_{i}}$ are linearly independent over the rational numbers, then
$\left|\beta_{1}\ln{\alpha_{1}}+\cdots+\beta_{n}\ln{\alpha_{n}}\right|>H^{-C},$
where $H$ is the maximum of the heights of the $\beta_{i}$ and $C$ is a function of $n$, the numbers $\alpha_{i}$, and the degrees of the numbers $\beta_{i}$.
Consider arithmetic expressions of the form $\beta_{1}\ln{\alpha_{1}}+\beta_{2}\ln{\alpha_{2}}$ where $\alpha_{1},\alpha_{2}\in\overline{\mathbb{Q}}\setminus 0$ are fixed and $\beta_{1},\beta_{2}\in\mathbb{Z}\setminus 0$. Now $\beta_{1}$ and $\beta_{2}$ have degree 111, so the exponent $C$ becomes a constant, and $H=\max\{|\beta_{1}|,|\beta_{2}|\}$. Applying Baker’s theorem (Theorem 2.4) and dividing both sides by $|\beta_{1}\ln{\alpha_{2}}|$ gives
$\left|\frac{\ln{\alpha_{1}}}{\ln{\alpha_{2}}}+\frac{\beta_{2}}{\beta_{1}}\right|>\frac{1}{|\beta_{1}\ln{\alpha_{2}}|H^{C}}\geq\frac{1}{|\ln{\alpha_{2}}|H^{C+1}}.$
By choosing $\alpha_{1}=\frac{2+i}{\sqrt{5}}$ and $\alpha_{2}=-1$, the ratio $\frac{\ln{\alpha_{1}}}{\ln{\alpha_{2}}}$ is exactly $\gamma/\pi$. Recall that this number is irrational, satisfying the criteria that $\ln{\alpha_{1}}$ and $\ln{\alpha_{2}}$ are linearly independent over the rationals. Since $\gamma/\pi<1$, clearly we can replace $H$ by $\beta_{1}$ to obtain $\left|\frac{\ln{\alpha_{1}}}{\ln{\alpha_{2}}}+\frac{\beta_{2}}{\beta_{1}}\right|>\frac{1}{|\ln{\alpha_{2}}|\beta_{1}^{C+1}}$. Comparing this with the definition of the irrationality exponent, this is sufficient to see that $\mu(\gamma/\pi)$ is at most $C+1$ and therefore finite.
Henceforth let $\mu=\mu(\gamma/\pi)$. It is worth noting that the constant $C$ given by Theorem 2.4 is effectively computable, though the order of magnitude is impractical. [^1] One estimate, also due to Baker [Bak76], states that $C=\ln{A_{1}}\ln^{2}{A_{2}}(16nd)^{200n}$ suffices when the $\beta_{i}$ are rational, where $A_{i}$ is the height of $\alpha_{i}$ and $d$ is the degree of the field extension $\mathbb{Q}[\alpha_{1},\alpha_{2}]/\mathbb{Q}$. For our purposes, $A_{1}=4$, $A_{2}=2$, $n=2$, and $d=4$, giving $C\approx 2^{2800}$. Baker remarks it is possible to argue $C$ is much lower in reality, but the required analysis is quite technical and beyond the scope of this paper.
2.3 Convergence of Stretch
We define $f(d)$ to be the maximum stretch guaranteed by $G_{\omega}$, over all pairs of vertices at Euclidean distance at least $d$.
$f(d)=\sup\left\{\frac{\operatorname{dist}_{G_{\omega}}(u,v)}{\|u-v\|_{2}}\;\middle|\;u,v\in V(G_{\omega})\text{ and }\|u-v\|_{2}\geq d\right\}.$
Lemma 2.1.
Fix any distance $d$ and let $f(d)=1+\epsilon$. If $d=\Omega(5^{\Theta(\varepsilon^{(1-\mu)/2})})$, then
$f(3d)<1+\epsilon-\Omega\left(5^{-\Theta(\varepsilon^{(1-\mu)/2})}\right).$
Proof.
Consider any two vertices $P,Q\in G_{\omega}$ with $|\overline{PQ}|=D\geq 3d$ when we regard them as points in the plane. Begin by choosing a parameter $\delta=\delta(\epsilon)$ and identifying a triangle $T$ satisfying the following properties: $T$ intersects $\overline{PQ}$, the projection of $T$ onto $\overline{PQ}$ lies entirely within the middle third of $\overline{PQ}$, and the hypotenuse of $T$ has length exactly $\delta D$. Now choose a parameter $\theta=\theta(\epsilon)$ and let $n=n(\theta)$ be large enough such that $\mathrm{Angles}(T,n)$ is a $\theta$-cover of $[0,\pi)$. Finally, select a triangle $T^{\prime}$ within $T$, and exactly $n$ levels below $T$, whose hypotenuse creates an angle less than $\theta$ with $\overline{PQ}$.
We now construct a path as follows. Let $p$ and $q$ be the endpoints of the hypotenuse of $T^{\prime}$ closer to $P$ and $Q$ respectively. Let $L_{1}=\overline{Pp}$, $L_{2}=\overline{pq}$, and $L_{3}=\overline{Qq}$. $|L_{1}|$ and $|L_{3}|$ must each be at least $d$, so by the definition of the stretch function $f$, $G_{\omega}$ approximates $L_{1}$ and $L_{3}$ to within $1+\varepsilon$ stretch.
Let $\ell_{1}$, $\ell_{2}$, and $\ell_{3}$ be the projections of $L_{1}$, $L_{2}$, and $L_{3}$ onto $\overline{PQ}$. We claim that $\operatorname{dist}_{G_{\omega}}(P,p)/|\ell_{1}|<(1+\varepsilon)(1+5\delta^{2})$. Since $p$ lies within $T$, its projection onto $\overline{PQ}$ is a distance at most $\delta D$ from $p$ and at least $D/3$ from $P$. Therefore, $|L_{1}|/|\ell_{1}|$ is bounded by $\sqrt{(D/3)^{2}+(\delta D)^{2}}/(D/3)=\sqrt{1+9\delta^{2}}<1+5\delta^{2}$. By definition of $f$, $\operatorname{dist}_{G_{\omega}}(P,p)/|L_{1}|\leq 1+\varepsilon$. The same is true for $\operatorname{dist}_{G_{\omega}}(q,Q)/|\ell_{3}|$. On the other hand, $\operatorname{dist}_{G_{\omega}}(p,q)/|\ell_{2}|$ is trivially bounded by $|L_{2}|/|\ell_{2}|\leq 1/\cos\theta$ since $p$ and $q$ are joined directly by an edge in $G_{\omega}$. Combining the bounds on each component, we write
$\operatorname{dist}_{G_{\omega}}(P,Q)<(1+\varepsilon)(1+5\delta^{2})(|\ell_{1}|+|\ell_{3}|)+\frac{|\ell_{2}|}{\cos\theta}.$
We normalize by $1/D$ to obtain the stretch, and because this inequality holds for arbitrary $P,Q$ satisfying $|\overline{PQ}|\geq 3d$, it holds for the supremum as well.
$f(3d)$ $=\sup_{P,Q\in V(G_{\omega})}\left\{\frac{\operatorname{dist}_{G_{\omega}}(P,Q)}{|\overline{PQ}|}\;\middle|\;|\overline{PQ}|\geq 3d\right\}$
$<(1+\varepsilon)(1+5\delta^{2})\left(1-\frac{|\ell_{2}|}{D}\right)+\frac{|\ell_{2}|}{D\cos\theta}$
$\leq(1+\varepsilon)(1+5\delta^{2})(1-5^{-n/2}\delta\cos\theta)+5^{-n/2}\delta$
Letting $\kappa=5^{-n/2}$, this is upper bounded by
$<1+\varepsilon+6\delta^{2}+\kappa\delta(1-(1+\varepsilon)\cos\theta)$
At this point we fix $\theta=\sqrt{\epsilon}$, so $\cos\theta>1-\epsilon/2$.
$<1+\varepsilon+6\delta^{2}-\kappa\delta\epsilon(1-\epsilon)/2=1+\varepsilon+\delta(6\delta-\kappa\epsilon(1-\epsilon)/2)$
Finally, we pick $\delta<\kappa\epsilon/24$.
$\leq 1+\varepsilon-\Omega(\delta^{2})=1+\epsilon-\Omega(5^{-n}\epsilon^{2})=1+\epsilon-\Omega(5^{\Theta(\epsilon^{(1-\mu)/2})}).$
The last line follows from Corollary 2.3, which states $n(\theta)=n(\sqrt{\epsilon})=(\sqrt{\epsilon})^{1-\mu}$.
The only remaining detail is to address that the triangles $T$ and $T^{\prime}$ used in this argument do indeed exist. The length of the hypotenuse of $T$ was taken to be $\delta D$, and then we chose $T^{\prime}$ to be $n$ levels beneath that of $T$. Therefore, we have implicitly assumed $D\geq 5^{n/2}/\delta>\Omega(5^{n}/\varepsilon)$, consistent with the assumption in the statement of the lemma. ∎
Theorem 2.1 now follows easily from this lemma.
Proof.
Take some initial constants $D_{0},\varepsilon_{0}$ with $f(D_{0})=1+\varepsilon_{0}$ and consider the sequence $\{\varepsilon_{i}\}$ where $\varepsilon_{i}=f(3^{i}D_{0})-1$. Define $i_{\operatorname{half}}$ to be the minimum value such that $\varepsilon_{i_{\operatorname{half}}}\leq\varepsilon_{0}/2$. By Lemma 2.1, if $D$ is sufficiently large, we have that $\varepsilon_{i}-\varepsilon_{i+1}>\Omega\left(5^{-\Theta\left(\varepsilon_{i}^{(1-\mu)/2}\right)}\right)>\Omega\left(5^{-\Theta\left(\varepsilon_{0}^{(1-\mu)/2}\right)}\right)$ as long as $\varepsilon_{i}>\varepsilon_{0}/2$. Therefore, $i_{\operatorname{half}}=O\left(5^{\Theta\left(\varepsilon_{0}^{(1-\mu)/2}\right)}\right)$. If our target stretch is $1+\varepsilon$ we can apply this halving argument $k=\log\epsilon^{-1}$ times, implying $f(3^{i^{*}}D_{0})<1+\varepsilon$ for
$i^{*}=O\left(5^{\Theta\left(\varepsilon_{0}^{(1-\mu)/2}\right)}+5^{\Theta\left((\varepsilon_{0}/2)^{(1-\mu)/2}\right)}+\cdots+5^{\Theta\left((\varepsilon_{0}/2^{k})^{(1-\mu)/2}\right)}\right)=O\left(5^{\Theta\left(\varepsilon^{(1-\mu)/2}\right)}\right).$
Note that for $D=3^{i^{*}}D_{0}$, $\epsilon=\Theta(\log^{\xi}\log D)$, for $\xi=2/(1-\mu)=\Theta(1)$. ∎
2.4 Pinwheel Tilings on the Grid
Theorem 2.5.
There exists a weight function $w$ such that $\mathsf{Grid}[w]$ has stretch 1, asymptotically. In particular, for any $u,v\in V(\mathsf{Grid})$,
$\|u-v\|_{2}-O(1)\leq\operatorname{dist}_{w}(u,v)\leq(1+O(1/(\log\log\|u-v\|_{2})^{\xi}))\cdot\|u-v\|_{2},$
for some $\xi>0$.
Proof.
Regard $G_{\omega}$ as the plane graph of a tiling whose atomic tile has large side lengths, say $25,50,25\sqrt{5}$. The vertices of $G_{\omega}$ generally do not have integer coordinates. Let $\phi:V(G_{\omega})\to V(\mathsf{Grid})$ map any $u\in V(G_{\omega})$ to the nearest integer point $\phi(u)\in\mathbb{Z}^{2}$. We will overload this notation a bit and let $\phi:E(G_{\omega})\to 2^{E(\mathsf{Grid})}$ be such that $\phi(\{u,v\})$ is a monotone path in $\mathsf{Grid}$ connecting $\phi(u)$ to $\phi(v)$ cleaving closely to the $\overline{uv}$ line segment, with the property that any two paths $\phi(\{u,v\}),\phi(\{u,v^{\prime}\})$ only intersect in a prefix of at most 2 edges, and any $\phi(\{u,v\}),\phi(\{u^{\prime},v^{\prime}\})$ (with $u,v,u^{\prime},v^{\prime}$ distinct) do not intersect at all.
The edge weights are assigned as follows. If $e$ is not in $\bigcup_{e^{\prime}\in E(G_{\omega})}\phi(e^{\prime})$ then $w(e)=10$. If $e$ is in two distinct paths $\phi(e^{\prime}),\phi(e^{\prime\prime})$ then $w(e)=1$. The remaining edge weights of $\phi(\{u,v\})$ are chosen to be equal, such that
$w(\phi(\{u,v\}))=\sum_{e\in\phi(\{u,v\})}w(e)=\|u-v\|_{2}.$
In other words, walking from $\phi(u)$ to $\phi(v)$ along $\phi(\{u,v\})$ is precisely the Euclidean distance $\|u-v\|_{2}$. Depending on the angle of the $u$-$v$ line, the “ideal” weight of edges on $\phi(\{u,v\})$ is in the range $[1/\sqrt{2},1]$, but the true weights lie in the range [0.6,1.05][0.6,1.05][ 0.6 , 1.05 ]. The internal edges of $\phi(\{u,v\})$ may need to have weight less than $1/\sqrt{2}$ due to rounding $u,v$ to farther integer points $\phi(u),\phi(v)$, and correcting for up to four edges on the ends of $\phi(\{u,v\})$ with weight 1. Similarly, the internal edges of $\phi(\{u,v\})$ may need to have weight greater than 1 due to rounding $u,v$ to closer integer points $\phi(u),\phi(v)$. However, one may verify that the length of every subpath of $\phi(\{u,v\})$ from $u^{\prime}$ to $v^{\prime}$ differs from its Euclidean length $\|u^{\prime}-v^{\prime}\|_{2}$ by at most 2.
By Theorem 2.1, for any $u,v\in V(\mathsf{Grid})$, $\operatorname{dist}_{w}(u,v)\leq(1+O(1/(\log\log\|u-v\|_{2})^{\xi}))\cdot\|u-v\|_{2}$. One walks from $u$ to a nearby $\phi(u_{0})$ vertex, then along embedded paths of the pinwheel graph $G_{\omega}$ to a $\phi(v_{0})$ near $v$, then along a path from $\phi(v_{0})$ to $v$. By design, the length of the path from $\phi(u_{0})$ to $\phi(v_{0})$ is precisely $\operatorname{dist}_{G_{\omega}}(u_{0},v_{0})$, while the $u$-$u_{0}$ and $v_{0}$-$v$ paths have length $O(1)$. The weight of edges outside of $\bigcup_{e\in E(G_{\omega})}\phi(e)$ is set sufficiently high so that it is never advantageous to use them in lieu of paths in $\bigcup_{e\in E(G_{\omega})}\phi(e)$. ∎
3 A Deterministic Construction with Faster Convergence
In this section, we give a deterministic construction based on highways with faster convergence. Specifically, we establish the following result as Theorem 3.1:
Theorem 3.1.
There exists an assignment $W:E(\mathsf{Grid})\to\mathbb{R}_{\geq 0}$ such that for any $u,v\in V(\mathsf{Grid})$,
$\|u-v\|_{2}-1\leq\operatorname{dist}_{W}(u,v)\leq\|u-v\|_{2}+O\left(\|u-v\|_{2}^{\frac{8}{9}}\right).$
We prove Theorem 3.1 in two steps. First, we show that the same statement holds for the finite square grid $[n]\times[n]$ (Theorem 3.2), then we give a black-box reduction from the infinite case to the finite case.
Theorem 3.2.
There exists a weight assignment $W(n)$ to edges of the finite grid on $[n]\times[n]$ such that for any $u,v\in[n]^{2}$,
$\|u-v\|_{2}-1\leq\operatorname{dist}_{W(n)}(u,v)\leq\|u-v\|_{2}+O\left(\|u-v\|_{2}^{\frac{8}{9}}\right).$
The proof of Theorem 3.2 follows from the constructions presented in Section 3.1 and Section 3.2.
3.1 Highways
Given a line $\ell=ax+b$ in $\mathbb{R}^{2}$ we define the $\operatorname{Highway}(\ell)$ to be a grid-path that tracks $\ell$; see Fig. 3.
Figure 3: A grid-path that tracks $\ell$.
Specifically, let $v\in V(\mathsf{Grid})$ be a grid-point, and consider the $\times 1$ 1 square $[v-0.5,v+0.5)\times[v-0.5,v+0.5)$. If $\ell$ intersects this square, then we include $v$ in $V(\ell)$. Whenever $u,v\in V(\ell)$ are adjacent grid points, $\operatorname{Highway}(\ell)$ contains the edge $\{u,v\}\in E(\mathsf{Grid})$. If $a\in[-1,1]$ we let $w_{\ell}$ assign every edge in $\operatorname{Highway}(\ell)$ the weight $\frac{\sqrt{a^{2}+1}}{|a|+1}$, which is the asymptotic ratio between the Euclidean distance of points on $\ell$ and the number of edges taken along the grid path $\operatorname{Highway}(\ell)$. Otherwise we write $\ell$ as $x=a^{-1}(y-b)$ with $a^{-1}\in[-1,1]$ and use weight $\frac{\sqrt{a^{-2}+1}}{|a^{-1}|+1}$. All off-highway edges have weight $\infty$.
This weight assignment guarantees a discrepancy of at most 1 between Euclidean distances and grid distances along $\operatorname{Highway}(\ell)$. The proof of Lemma 3.1 appears in the appendix.
Lemma 3.1.
Let $\operatorname{Highway}(\ell)$ denote the highway that approximates a line $\ell$ of the form $y=ax+b$. Then for any two points $u,v\in V(\operatorname{Highway}(\ell))$, $\left|\operatorname{dist}_{w_{\ell}}(u,v)-\|u-v\|_{2}\right|\leq 1$.
The highway transformation can also be applied to a line segment $s$. We use the same notation $\operatorname{Highway}(s)$.
3.2 The Hierarchical Highway Construction
We are trying to find a weight assignment for the finite grid $[n]\times[n]$ in order to prove Theorem 3.2.
Parameters.
The construction is parameterized by $(k_{i})_{1\leq i\leq m}$, where
$k_{1}=\lfloor n^{1/5}\rfloor,\quad k_{i+1}=\lfloor k_{i}^{1/2}\rfloor,$
and $m$ is minimum such that $k_{m}<100$.
Layers of Lines.
The construction is based on a hierarchical system of lines in $\mathbb{R}^{2}$ which will eventually be embedded as highways in the grid. The lines at level $i$ have angles selected from $(\theta_{i,j})_{0\leq j<k_{i}}$:
$\theta_{i,j}=\frac{\pi\cdot j}{k_{i}}.$
Fixing $i$ and one such angle $\theta_{i,j}$, there are many lines with angle $\theta_{i,j}$, spaced at distance $k_{i}^{4}$. For $i\in[1,m],j\in[0,k_{i}-1],t\in\mathbb{Z}$,
$\ell_{i,j,t}=\left\{(x,y)\in\mathbb{R}^{2}\ \middle|\ y\cdot\cos\theta=x\cdot\sin\theta+t\cdot k_{i}^{4}\right\}.$
Define $\mathsf{Lines}[i]=\{\ell_{i,j,t}\}$ to be the set of all lines at level $i$.
We cannot choose a weight function $W(n)$ that agrees with $w_{\ell_{i,j,t}}$ for every line $\ell_{i,j,t}$ due to intersections. Our solution is to avoid this issue by removing all line intersections, which introduces distortions in distances that must be bounded.
Below we define a procedure to remove parts of $\ell_{i,j,t}$, leaving a set of line segments $\mathcal{L}_{i,j,t}$. Define $\mathsf{Segments}[i]=\bigcup\mathcal{L}_{i,j,t}$ to be the set of all line segments at level $i$. If $O$ is an object or collection of objects, define
$\operatorname{Fat}(O,\delta)=\{p\mid\exists q\in O\mbox{ such that }\|p-q\|_{2}\leq\delta\}$
to be all points within distance $\delta$ of $O$. Specifically, $\operatorname{Fat}(\ell,\delta)$ is a strip if $\ell$ is a line, and a hippodrome if $\ell$ is a segment.
Once $\mathsf{Segments}[1],\ldots,\mathsf{Segments}[i-1]$ are constructed, we construct $\mathsf{Segments}[i]$ as follows. For each $\ell_{i,j,t}\in\mathsf{Lines}[i]$ initialize $\ell\leftarrow\ell_{i,j,t}$ and proceed to remove parts of $\ell$ in Steps 1 and 2.
Step 1.
Set $\ell\leftarrow\ell-\operatorname{Fat}(\mathsf{Lines}[i]-\{\ell\},k_{i})$, i.e., we remove every part of $\ell$ within distance $k_{i}$ of any other line at level $i$.
Step 2.
For each segment $s\in\bigcup_{i^{\prime}<i}\mathsf{Segments}[i^{\prime}]$ such that $\operatorname{Fat}(s,k_{i})\cap\ell\neq\emptyset$, let $\overline{AB}=\operatorname{Fat}(s,k_{i})\cap\ell$. If $\|A-B\|_{2}\geq k_{i}$, set $\ell\leftarrow\ell-\overline{AB}$. Otherwise, let $B^{\prime}\in\ell_{i,j,t}$ be such that $\|A-B^{\prime}\|=k_{i}$ and $B\in\overline{AB^{\prime}}$, and set $\ell\leftarrow\ell-\overline{AB^{\prime}}$. See Fig. 4. Define $\mathcal{L}_{i,j,t}=\ell$, and include all line segments of $\mathcal{L}_{i,j,t}$ in $\mathsf{Segments}[i]$.
Figure 4: Illustrations of various cases in Step 2. (A) $\ell$ and $s$ intersect. The segment $\ell\cap\operatorname{Fat}(s,k_{i})$ is removed from $\ell$. (B) $\ell$ and $s$ do not intersect. Left: $\overline{AB}\geq k_{i}$ and $\overline{AB}$ is removed from $\ell$. Right: $\overline{AB}<k_{i}$ and $B^{\prime}$ is such that $\overline{AB^{\prime}}=k_{i}$ is removed from $\ell$.
The weight assignment $W(n)$ is now constructed as follows. For each line segment $s\in\bigcup_{i}\mathsf{Segments}[i]$, let $W(n)$ agree with $w_{s}$ at all edges in the corresponding highway segment $\operatorname{Highway}(s)$. All edges not appearing in any line segment have weight 2.
Lemma 3.2.
Fix any $s\in\mathsf{Segments}[i]$, $s^{\prime}\in\mathsf{Segments}[i^{\prime}]$, where $s\neq s^{\prime}$ and $i\geq i^{\prime}$. For any points $u\in s,u^{\prime}\in s^{\prime}$, $\|u-v\|_{2}\geq k_{i}$.
Proof.
Steps 1 and 2 ensure that all remaining points on $\mathcal{L}_{i,j,t}$ lie outside
$\operatorname{Fat}\left(\bigcup_{i^{\prime}<i}\mathsf{Segments}[i^{\prime}]\cup(\mathsf{Lines}[i]-\{\ell_{i,j,t}\}),k_{i}\right),$
and that any two consecutive segments on $\mathcal{L}_{i,j,t}$ are separated by distance at least $k_{i}$. This implies that the Euclidean distance between any $u\in s$ and $u^{\prime}\in s^{\prime}$ is at least $k_{i}$. ∎
Lemma 3.3 shows that distances under $W(n)$ are approximately Euclidean, up to a multiplicative stretch of $1+O(k_{i}^{-1})$ and additive stretch $O(k_{i}^{4})$, for every index $i$.
Lemma 3.3.
For any $u,v\in[n]^{2}$ and $i\in[m]$.
$\|u-v\|_{2}-1\leq\operatorname{dist}_{W(n)}(u,v)\leq\|u-v\|_{2}+O\left(k_{i}^{-1}\|u-v\|_{2}+k_{i}^{4}\right).$
Proof.
Observe that the highways corresponding to all segments in $\bigcup_{i}\mathsf{Segments}[i]$ are vertex-disjoint. Thus, every grid path $P$ from $u$ to $v$ can be written as $B_{0}A_{1}B_{1}A_{2}B_{2}\cdots A_{k}B_{k}$, where each $A_{j}\subset\operatorname{Highway}(s)$ for some $s\in\bigcup_{i}\mathsf{Segments}[i]$ and each $B_{j}$ is disjoint from all highway segments, and therefore consists of only weight-2 edges. (One or both of $B_{0},B_{k}$ may be empty.) By Lemma 3.1, $W(A_{j})$ is at least the Euclidean distance between its endpoints minus one, and $W(B_{j})$ is at least the Euclidean distance multiplied by 2, which implies that $W(B_{j})$ is at least the Euclidean distance plus 1 for $1\leq j\leq k-1$, so $W(P)$ is at least $\|u-v\|_{2}-1$.
[…]
$\operatorname{dist}_{W(n)}(u,A)+\operatorname{dist}_{W(n)}(v,B)=O\left(k_{i}^{4}+\|u,v\|_{2}\sin\left(\frac{\pi}{2k_{i}}\right)\right)=O(k_{i}^{4}+k_{i}^{-1}\|u-v\|_{2}).$ (2)
It remains to bound $\operatorname{dist}_{W(n)}(A,B)$. A trivial upper bound is $\operatorname{dist}_{W(n)}(A,B)\leq 2\cdot\sqrt{2}\|A-B\|_{2}$, so if $\|A-B\|_{2}<100k_{i}^{4}$ we are done. Henceforth we shall assume that $\|A-B\|_{2}\geq 100k_{i}^{4}$. We would prefer to follow the $A$-$B$ path along $\operatorname{Highway}(\ell^{*})$, but sections of this highway have effectively been removed by Steps 1 and 2 of the construction. We bound the stretch induced by the gaps in the highway introduced in Steps 1 and 2 separately.
Step 1 Stretch.
Fix a direction $\theta_{i,j^{\prime}}$ different from $\ell^{*}$’s direction $\theta_{i,j}$. Whenever $\ell^{*}$ intersects a line with angle $\theta_{i,j^{\prime}}$, Step 1 causes $\operatorname{Highway}(\ell^{*})$ to lose $\frac{k_{i}}{\sin|\theta_{i,j}-\theta_{i,j^{\prime}}|}=O(k_{i}^{2})$ edges. There are $k_{i}$ angles, and parallel lines with angle $\theta_{i,j^{\prime}}$ are spaced $k_{i}^{4}$ apart, so the total number of edges removed from the $A$-$B$ path in $\operatorname{Highway}(\ell^{*})$ in Step 1 is
$O\left(k_{i}^{2}\cdot k_{i}\cdot\frac{\|A-B\|_{2}}{k_{i}^{4}}\right)=O(k_{i}^{-1}\|A-B\|_{2}).$ (3)
The additive stretch induced by walking across the gaps induced by Step 1 is also $O(k_{i}^{-1}\|A-B\|_{2})$ as all these edges have weight 2.
Step 2 Stretch.
Whenever part of $\operatorname{Highway}(\ell^{*})$ is removed by Step 2 we do not walk precisely in the direction of $\ell^{*}$ but take a detour to a lower level highway. Suppose that in Step 2, a segment $s\in\bigcup_{i^{\prime}<i}\mathsf{Segments}[i^{\prime}]$ causes an interval $\overline{CD}$ of $\ell^{*}$ to be removed. Define $E,F$ to be the points on $s$ closest to $C,D$, respectively. When our path reaches $C$, we walk from $C$ to $E$, then to $F$ along $\operatorname{Highway}(s)$, then to $D$. See Fig. 5. By construction $\|C-E\|_{2},\|D-F\|_{2}=O(k_{i})$. Since $\|E-F\|_{2}\leq\|C-D\|_{2}$ and by Lemma 3.1 $\operatorname{dist}_{W(n)}(E,F)=\|E-F\|_{2}\pm 1$, the additive stretch due to the conflict with $s$ is at most
$\operatorname{dist}_{W(n)}(C,E)+\operatorname{dist}_{W(n)}(E,F)+\operatorname{dist}_{W(n)}(F,D)-\operatorname{dist}_{w_{\ell^{*}}}(C,D)=O(k_{i}).$
The last task is to bound the number of such segments interfering with the $A$-$B$ path. Observe that $s$ is a segment of a line at level $i-1$ or lower. Thus, by Lemma 3.2 any two such segments $s,s^{\prime}$ are at distance at least $k_{i-1}\geq k_{i}^{2}$, and the total additive stretch caused by Step 2 detours is
$O\left(k_{i}\cdot\frac{\|A-B\|_{2}}{k_{i}^{2}}\right)=O(k_{i}^{-1}\|A-B\|_{2}).$ (4)
Figure 5: Step 2 detours from the proof of Lemma 3.3. (A) The case when segment $s$ intersects $\ell^{*}$. (B) When segment $s$ does not intersect $\ell^{*}$, but $\operatorname{Fat}(s,k_{i})$ does.
Combining Eqs. 2, 3 and 4, we conclude that
$\operatorname{dist}_{W(n)}(u,v)$ $\leq\operatorname{dist}_{W(n)}(u,A)+\operatorname{dist}_{W(n)}(A,B)+\operatorname{dist}_{W(n)}(B,v)$
$\leq\|u-v\|+O\left(k^{-1}\|A-B\|_{2}+k_{i}^{-1}\|u-v\|_{2}+k_{i}^{4}\right)$
$=\|u-v\|_{2}+O\left(k^{-1}\|u-v\|_{2}+k_{k}^{4}\right).$
Proof of Theorem 3.2. Recall that by the definition of the sequence $(k_{i})$, for any pair of points $u,v$ with Euclidean distance $d=\|u-v\|_{2}>100^{9}$, there exists some index $1\leq i\leq m$ such that
$k_{i}\in\left[d^{1/9}-1,\ d^{2/9}\right].$
Applying Lemma 3.3, we have
$\operatorname{dist}_{W}(u,v)$ $\leq\|u-v\|_{2}+O\left(\frac{1}{k_{i}}\cdot\|u-v\|_{2}+k_{i}^{4}\right)$
$\leq\|u-v\|_{2}+O\left(\|u-v\|_{2}^{8/9}\right).$
The theorem holds trivially when $d\leq 100^{9}$, which completes the proof of Theorem 3.2. $\square$
To prove Theorem 3.1 we give a “black box” reduction showing that any construction that gives a bound like Theorem 3.2 for the finite grid $[n]\times[n]$ yields the same guarantee on the infinite grid $\mathbb{Z}\times\mathbb{Z}$.
We begin by tiling the integer grid with various size squares as follows. The central tile is $\times 1000$ 1000, which is surrounded by eight $\times 1000$ 1000 tiles, all of which, in turn, are surrounded by eight $\times 3000$ 3000 tiles, which are in turn surrounded by eight $\times 9000$ 9000 tiles, and so on. See Fig. 6. Within each of these $n\times n$ tiles, we apply Theorem 3.2 to choose the weight function in the central $(n-2)\times(n-2)$ grid. All edges with at least one endpoint on the boundary of the tile have weight 2. Let $W$ be the resulting weight function of $E(\mathsf{Grid})$.
Figure 6: Illustration of the recursive tiling: at each level, we place eight squares around the current one to expand the scale.
Proof of Theorem 3.1. Consider any two $u,v\in V(\mathsf{Grid})$ and let $L\geq 0$ be minimum such that $\|u-v\|_{2}\leq 3^{L}1000$. The line $\overline{uv}$ can intersect at most 3 tiles with dimensions $3^{L}1000$ or larger, and at most 4 tiles with dimension $3^{i}1000$, $i<L$. Thus, by concatenating shortest paths inside each tile, by Theorem 3.2, the total additive stretch is at most
$\operatorname{dist}_{W}(u,v)-\|u-v\|_{2}=O\left(\|u-v\|_{2}^{8/9}+\sum_{i=0}^{L-1}(\sqrt{2}\cdot 3^{i}1000)^{8/9}\right)=O(\|u-v\|_{2}^{8/9}).$
The same argument from Lemma 3.1 shows that $\operatorname{dist}_{W}(u,v)\geq\|u-v\|_{2}-1$. $\square$
4 Experimental Findings in the Squishy Grid
If Question 2 seems too daunting, a natural idea is to simplify the problem by considering only monotone paths, that is, paths that use the fewest number of edges.
Question 3.
Define $\operatorname{monodist}_{w}(u,v)$ to be the length of the shortest $u$-$v$ path that uses $\|u-v\|_{1}$ edges. Does there exist a distribution $\mathscr{D}^{*}$ over $\mathbb{R}_{\geq 0}$ such that if $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}^{*}]$ is a randomly weighted graph, for all $u,v\in V(\mathsf{Grid})$,
${\mathbb{E}\/}(\operatorname{monodist}_{w}(u,v))=(1\pm o(1))\|u-v\|_{2}.$
At first glance this problem may seem easier, or more plausible, than Question 2. Whereas it is an open problem finding a distribution $\mathscr{D}$ satisfying Eq. 1 (the time constant in the $0^{\circ}$ and $45^{\circ}$ directions are 1), this is nearly trivial when we consider $\operatorname{monodist}$.
Lemma 4.1.
There exists a distribution $\mathscr{D}$ on $\mathbb{R}_{\geq 0}$ such that
$\lim_{n\to\infty}\frac{{\mathbb{E}\/}(\operatorname{monodist}_{w}(\mathbf{0},ne_{0}))}{n}=\lim_{n\to\infty}\frac{{\mathbb{E}\/}(\operatorname{monodist}_{w}(\mathbf{0},ne_{45}))}{n}=1.$
Proof.
Since there is only one path from $\mathbf{0}$ to $ne_{0}$, any distribution $\mathscr{D}$ with ${\mathbb{E}\/}(w_{0}\sim\mathscr{D})=1$ works for the $0^{\circ}$ direction. Consider the class of distributions $\mathscr{D}[\epsilon]$, where $\Pr(w_{0}=1-\epsilon)=\Pr(w_{0}=1+\epsilon)=1/2$. When $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}[0]]$, $\operatorname{monodist}_{w}(\mathbf{0},ne_{45})=\sqrt{2}\cdot\|u-v\|_{2}$. We argue that when $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}[1]]$, ${\mathbb{E}\/}[\operatorname{monodist}_{w}(\mathbf{0},ne_{45})]<\|u-v\|_{2}/\sqrt{2}+O(\sqrt{n})$. When $\epsilon=1$ all weights are 0 or 2 with equal probability. We walk myopically from the origin, taking a weight-0 edge North or East whenever possible, or a weight-2 edge North or East if necessary, until we reach a barrier when the $x$- or $y$-coordinate matches $ne_{45}$. When the edges in both directions have the same weight, we choose one randomly. Before reaching a barrier, the expected weight of the next edge is (3/4)0+(1/4)2=1/2(3/4$\cdot$ 0+(1/4$\cdot$ 2=1/2( 3 / 4 ) ⋅ 0 + ( 1 / 4 ) ⋅ 2 = 1 / 2 and after reaching a barrier it is (1/2)(0+2)=1(1/2)(0+2)=1( 1 / 2 ) ( 0 + 2 ) = 1. There are $O(\sqrt{n})$ edges in the latter category, in expectation, so ${\mathbb{E}\/}(\operatorname{monodist}_{w}(\mathbf{0},ne_{45}))<(1/2)\|ne_{45}\|_{1}+O(\sqrt{n})=n/\sqrt{2}+O(\sqrt{n})$. By the intermediate value theorem, there has to be some $\epsilon^{*}\in[0,1]$ such that
$\lim_{n\to\infty}\frac{{\mathbb{E}\/}(\operatorname{monodist}_{w}(\mathbf{0},ne_{0}))}{n}=\lim_{n\to\infty}\frac{{\mathbb{E}\/}(\operatorname{monodist}_{w}(\mathbf{0},ne_{45}))}{n}=1.$
Let $B_{\operatorname{mono}}(t)=\{u\in\mathbb{Z}^{2}\mid\operatorname{monodist}_{w}(\mathbf{0},u)\leq t\}$ and $\mathcal{B}_{\operatorname{mono}}(\mathscr{D})$ be the limiting shape of $B_{\operatorname{mono}}(t)/t$ in $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}]$ as $t\to\infty$. Thus, $\mathcal{B}_{\operatorname{mono}}(\mathscr{D}_{\epsilon^{*}})$ coincides with the $L_{2}$-ball in the eight intercardinal directions N, E, S, W, NE, SE, SW, NW. If $\mathcal{B}_{\operatorname{mono}}(\mathscr{D}_{\epsilon^{*}})$ were convex, then it would have to be quite close to the $L_{2}$-ball.
Unfortunately, our experiments show that $\mathcal{B}_{\operatorname{mono}}(\mathscr{D}_{\epsilon^{*}})$ is not convex, which casts serious doubt on Question 3 having an affirmative answer. See Fig. 7.
Refer to caption Refer to caption
(A) (B)
Figure 7: (A) The stretch of $\operatorname{monodist}_{w}(\mathbf{0},ne_{\theta})/\|ne_{\theta}\|_{2}$, as a function of the angle $\theta\in[0,\frac{\pi}{2}]$, expressed in radians. (B) The shape of $\mathcal{B}_{\operatorname{mono}}(\mathscr{D}_{\epsilon^{*}})$.
In retrospect, Question 3 is less likely than Question 2 to be answered in the affirmative since $\operatorname{monodist}_{w}(\mathbf{0},(n,m))$, $m\leq n$, is much more sensitive to small deviations in $m$ than $\operatorname{dist}_{w}(\mathbf{0},(n,m))$ Considering the cases when $m=0,m=0.1n$, and $m=n$, $\operatorname{monodist}_{w}(\mathbf{0},(n,m))$ is the minimum of (${n+0\choose 0}=1,{n+0.1n\choose 0.1n}\approx 1.39^{n}$, and (${2n\choose n}\approx 4^{n}$ different paths, respectively. This sharp jump from constant to exponential in the vicinity $m=0$ does not exist in Question 2. Assuming the variance of $\mathscr{D}$ is sufficiently large, $\operatorname{dist}_{w}(\mathbf{0},ne_{0})$ is the minimum of an exponential number of plausible shortest paths.
4.1 Discrete Distributions for First Passage Percolation
When dealing with discrete distributions the most natural measure of complexity is support size. Therefore, we study Question 2 experimentally by considering the space of 2- and 3-point distributions. For a fixed integer $k$, the $k$-point distribution $\mathscr{D}(\{(p_{i},x_{i})\}_{i=1}^{k})$, is such that
$\Pr_{w_{0}\sim\mathscr{D}}(w_{0}=x_{i})=p_{i}.$
It is determined by $2k-1$ parameters, as $p_{k}=1-(p_{1}+\cdots+p_{k-1})$.
4.1.1 Experimental Methodology
To identify locally optimal distributions in the space of $k$-point discrete distribution $\mathscr{D}(\{(p_{i},x_{i})\}_{i=1}^{k})$, we employ a two-layer iterative strategy:
  • We first perturb the probability vector $(p_{1},\ldots,p_{k})$.
  • For fixed probabilities $(p_{1},\ldots,p_{k})$, we generate $k$ random initial values $x_{1},x_{2},\ldots,x_{k}$ and then alternate between the following two update steps:
    Perturbation Step.
    We perturb each value $x_{i}$ and compute the estimated directional stretch for both $\theta=0$ and $\theta=\frac{\pi}{4}$. Let $\mu_{\theta}$ be the empirical ratio $\frac{\operatorname{dist}_{w}(\mathbf{0},ne_{\theta})}{n}$, obtained from this round of simulation. Here $n\approx$ 30,000.
    Normalization Step.
    We normalize the values $\{x_{i}\}_{i=1}^{k}$ by setting
    $x_{i}\leftarrow\frac{x_{i}}{\sqrt{\mu_{0}\cdot\mu_{\frac{\pi}{4}}}}.$
    This scaling ensures that the average stretch along the cardinal and intercardinal directions remains close to 1, thereby facilitating comparisons between distributions.
4.1.2 2-Point Distributions
The best 222-point distribution identified with this method is $\mathscr{D}_{2}$, given below. Roughly speaking, every edge weight is either 0.410.410.41 or 4.754.754.75, 444444% and 565656% of the time, respectively.
$\mathscr{D}_{2}=\{(0.44273,0.41401),(0.55727,4.75309)\}.$
[…]
Refer to caption Refer to caption
(A) (B)
[…]
Note that ${\mathbb{E}\/}(w_{0}\sim\mathscr{D}_{2})\approx 2.83$, meaning that $\operatorname{dist}_{w}(\mathbf{0},ne_{0})\approx n$ is likely to be realized by a highly non-monotone path, consistent with the observations in Fig. 7. Two sample paths from the origin to (1000,0)(1000,0)( 1000 , 0 ) and (1000,100)(1000,100)( 1000 , 100 ) are shown in Fig. 9.
Refer to caption Refer to caption
(A) (B)
Figure 9: Results on the 2-point distribution $\mathscr{D}_{2}$. (A) The trace of a shortest path from (0,0)(0,0)( 0 , 0 ) to (1000,0)(1000,0)( 1000 , 0 ). (B) The trace of a shortest path from (0,0)(0,0)( 0 , 0 ) to (1000,100)(1000,100)( 1000 , 100 ).
4.1.3 3-Point Distributions
$\mathscr{D}_{2}$ does not leave much room for improvement, but we are able to eke out a slightly better empirical stretch of 1.00622 with a 3-point distribution $\mathscr{D}_{3}$.
$\mathscr{D}_{3}$ $=\mathscr{D}(\{(0.34809,0.20647),(0.25735,2.51586),(0.39456,9.32215)\}).$
See Fig. 10 for visual representations of the empirical stretch of $\mathscr{D}_{3}$.
Refer to caption Refer to caption
(A) (B)
[…]
4.2 $L_{p}$-Balls and Continuous Distributions
Alm and Deijfen [AD15] experimented with many of the standard continuous distributions, such as uniform, exponential, Gamma, and Fisher. Only a distribution from the Fisher class approximated Euclidean distances to within 1%. In this section we replicate some of Alm and Deijfen’s findings, but instead of measuring error with respect to the $L_{2}$-norm, we show they are very good approximations for other $L_{p}$-norms, $p<2$. Fig. 11 shows that (suitable scaled versions of) $\mathrm{Uniform}(0,1)$, $\Gamma(2,2)$ and $\Gamma(10,10)$ are good approximations to the $L_{1.87},L_{1.85},$ and $L_{1.32}$ metrics, respectively. It is not true that every $\mathcal{B}(\mathscr{D})$ approximates an $L_{p}$-ball. For some non-constant distributions, the limit shape $\mathcal{B}(\mathscr{D})$ has flat edges; see [ADH17, §2.5].
Refer to caption Refer to caption Refer to caption
(A) (B) (C)
Figure 11: Other $L_{p}$ balls, $p\in(1,2)$, and rescaled versions of some continuous distributions. (A) $L_{1.87361}$ ball and $\mathrm{Uniform}(0,1)$ (B) $L_{1.85691}$ ball and $\Gamma(2,2)$ (C) $L_{1.32879}$ ball and $\Gamma(10,10)$
We measure the empirical error of the graph distances in $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}]$ with respect to an $L_{p}$-norm as follows.
$\operatorname{Err}(p;w)=\max_{\textbf{u}\in\operatorname{Ball}_{n}}\left\lvert\frac{\lVert\textbf{u}\rVert_{2}}{\lVert\textbf{u}_{0}\rVert_{2}}-\left(|\cos(\theta(\textbf{u}))|^{p}+|\sin(\theta(\textbf{u}))|^{p}\right)^{-\frac{1}{p}}\right\rvert.$
Here $n$ is a large constant, $\operatorname{Ball}_{n}$ is the set of all u such that $\operatorname{dist}_{w}(\mathbf{0},\textbf{u})\geq n$ but the immediate predecessor of u is at distance less than $n$, and $\textbf{u}_{0}$ is the first vertex of $\operatorname{Ball}_{n}$ on the $x$-axis. The angle of u is $\theta(\textbf{u})$. Intuitively, if the time constant of $\mathscr{D}$ is not 1, we can effectively make it 1 for these calculations by normalizing by $\|\textbf{u}_{0}\|$. Then $\operatorname{Err}(p;w)$ measures the maximum difference between the normalized distance to a $\textbf{u}\in\operatorname{Ball}_{n}$ and its idealized radial profile in the $L_{p}$ ball, evaluated in direction $\theta(\textbf{u})$.
Our experimental results show that for some continuous distributions $\mathscr{D}$, $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}]$ closely approximates an $L_{p}$-metric for $p<2$.
Distribution $\mathscr{D}$ $p$ value $\operatorname{Err}(p;w)$
$\mathrm{Uniform}(0,1)$ 1.87361 0.00462
$\Gamma(2,2)$ 1.85691 0.00986
$\Gamma(10,10)$ 1.32879 0.03658
5 Conclusion
In this paper we asked how well the integer grid graph $\mathsf{Grid}$ can approximate Euclidean distances, if weighted appropriately. We gave two deterministic weighting schemes that answer Question 1, the best one achieving a polynomial additive stretch, that is, ${\mathbb{E}\/}[\operatorname{dist}_{w}(u,v)]=\|u-v\|_{2}+O(\|u-v\|_{2}^{1-\delta})$ for $\delta=1/9$. Improving the additive stretch to something subpolynomial $\|u-v\|_{2}^{o(1)}$ seems to require a new approach to the problem. Our “highway” method seems incapable of achieving subpolynomial additive error, and even if Question 2 is answered affirmatively (choosing weights i.i.d. from some distribution $\mathscr{D}^{*}$), the tail bounds here are at least polynomial [ADH17].
We conjecture that no weighting achieves subpolynomial additive error.
Conjecture 5.1.
The additive error $\operatorname{dist}_{w}(u,v)-\|u-v\|_{2}$ of $w:E(\mathsf{Grid})\to\mathbb{R}_{\geq 0}$ is a function of $d=\|u-v\|_{2}$.
Weak Conjecture.
There is no $w$ with constant additive error $O(1)$, independent of $d$.
Strong Conjecture.
There is no $w$ with subpolynomial additive error $d^{o(1)}$.
[…]
It is an interesting open problem to prove that in $\mathsf{Grid}[w]\sim\mathsf{Grid}[\mathscr{D}_{2}]$, the expected error of $\operatorname{dist}_{w}$ is at most 1%, that is:
${\mathbb{E}\/}(\operatorname{dist}_{w}(u,v))=(1.005\pm 0.005\pm o(1))\|u-v\|_{2}.$
Acknowledgments.
We would like to thank Greg Bodwin for posing the problem to us, Gábor Tardos for pointing us to references [PPS90, RS96], and Wesley Pegden for suggesting we look into the percolation literature.
Appendix A Proofs
A.1 Proof of Lemma 3.1
Proof of Lemma 3.1.
[…]
$\left|\operatorname{dist}_{w_{\ell}}(u,v)-\|u-v\|_{2}\right|$ $=\left|(x+y)\frac{\sqrt{a^{2}+1}}{a+1}-\sqrt{x^{2}+y^{2}}\right|$
$=\left|\frac{(x+y)\sqrt{a^{2}+1}-\sqrt{x^{2}+y^{2}}(a+1)}{a+1}\right|$
$=\left|\frac{\left((x+y)\sqrt{a^{2}+1}\right)^{2}-\left(\sqrt{x^{2}+y^{2}}(a+1)\right)^{2}}{(a+1)[(x+y)\sqrt{a^{2}+1}+(a+1)\sqrt{x^{2}+y^{2}}]}\right|$
$=\left|\frac{(x+y)^{2}(a^{2}+1)-(a+1)^{2}(x^{2}+y^{2})}{(a+1)\left[(x+y)\sqrt{a^{2}+1}+(a+1)\sqrt{x^{2}+y^{2}}\right]}\right|$
$=\left|\frac{2(ax-y)(ay-x)}{(a+1)\left[(x+y)\sqrt{a^{2}+1}+(a+1)\sqrt{x^{2}+y^{2}}\right]}\right|$
$\leq\frac{2(1+a)x}{(1+a)[x+x]}$
=1=1.= 1 .